The expected return on the market is 15 percent with a standard deviation of 12.5 percent and the risk-free rate is 5 percentplease show workings and dont have to show in excel
Answer #1

Answer : C: 2 and 3 Only

The capital market line expressed by the following formula:

Rp = R; +(Rm – Rp) *oplom

Where

Rp is portfolio expected/required rate of return

Rf is risk free rate of return

Rm is market return

Sigma P is portfolio standard deviation

Sigma m is market standard deviation

Using the above formula for all four portfolio we get the below mentioned results:

Rf = 5%    Rm = 15%       Market Standard Deviation= 12.5%  

Portfolio 1:

Rp1 = 5 + (15-5)*28.75/12.5 = 28% Does not match with the given return i.e. 3.%

Rp2 = 5 + (15-5) * 25/12.5 = 25% Matches with the given expected return

Rp3 = 5 + (15-5)*6.25/12.5 = 10% Matches with the given expected return

Rp3 = 5 + (15-5) *5.5/12.5 = 9.4% Does not match with given expected return i.e. 9%

Therefore, the answer will be 2 and 3 only.